a radio beacon is located at (-1, 0), and another at (2, -1). A navigator's equipment tells her that the line joining her position to the first beacon has slope -5 and the line joining her positon to the second beacon has slope 3. What is the navigator's approximate location?
Answer by stanbon(75887)

a radio beacon is located at (-1, 0), and another at (2, -1). A navigator's equipment tells her that the line joining her position to the first beacon has slope -5 and the line joining her positon to the second beacon has slope 3. What is the navigator's approximate location?
---------------
Plot the two points.
Find the equation of the line thru (-1,0) with a slope of -5
y=mx+b
0=-5(-1)+b
b=-5
EQUATION #1: y=-5x-5
-------------
Find the equation of the line thru (2,-1) with a slope of 3
y=mx+b
-1=3*2+b
b=-7
EQUATION #2: y=3x-7
----------------------
Find the intersection of the two lines:
-5x-5=3x-7
2x=2
x=1
Substitute to find "y":
y=-5*1-5
y=-10
--------------
The point you want is (1,-10)
Cheers,
Stan H.